What value will make the equation true? Logx 81= -4
A) x= 1/9

B) x= 1/3

C) x=3

D) x=9
1
Expert's answer
2014-06-04T13:28:07-0400
By definition of logarithm, the base of logarithm is a positive real number not equal to 1. The equality logx81=-4 is equivalent to 81=x(-4), that is 1/x4=81, or x4=1/81, x4-1/81=0. Seek a positive x, factor the left-hand side (x2-1/9)(x2+1/9)=0,
(x-1/3)(x+1/3)(x2+1/9)=0, x>0. The answer is x=1/3.